OpenStudy (anonymous):

write 15% as a fraction in simplest form

OpenStudy (anonymous):

15/100 =3/20

OpenStudy (anonymous):

15% means "15 per 100", or written as a fraction:\[\frac{15}{100}\]This fraction can be reduced though, by dividing both the top and bottom by 5:\[\frac{15 \times \frac{1}{5}}{100 \times \frac{1}{5}} = \frac{3}{20}\]